OpenStudy (anonymous):

Evaluate the factorial expression: 24!20!

hartnn (hartnn):

you mean \(\dfrac{24!}{20!}\) ?

geerky42 (geerky42):

Question is \(\dfrac{24!}{20!}\) You are thinking about \(\dfrac{24}{20}\)